OpenStudy (anonymous):

6*0.222*7.0 Express your answer numerically using the appropriate number of significant figures.

OpenStudy (anonymous):

the answer is 9

OpenStudy (mathmale):

I see you have 3 different multiplicands: 6, 0.222, and 7.0. 6 is expressed to the nearest integer. The others are expressed to 3 and 1 decimal place accuracy, respectively. Your final result, the product of these 3 numbers, cannot be any more accurate than the least accurate of the multiplicands. Thus, we round off the product, 9.324, to 9 with NO decimal points.

OpenStudy (tkhunny):

That doesn't quote make sense. The ONLY way to get 9.324 is to retain ALL decimal places in ALL intermediate results - entirely ignoring Significant Figures considerations. Multiply in various orders (6*0.222)*7.0 = 1*(7.0) = 7 (6*7.0)*0.222 = 40*.222 = 9 <== Sort of hard to tell that "40" is only one significant digit. Writing as 4E1 would be more obvious. 6*(0.222*7.0) = 6*1.6 = 10 <== This one retains the most precision for the longest time. Maximum (Ceiling) Calculation 6.5*0.2225*7.05 = 10.1960625 Minimum (Floor) Calculation 5.5*0.2215*6.95 = 8.4668375 There are various ways to go about it. You should follow the instructions from your class. It's a lot more trouble to worry about intermediate results, but it may be the requirement. Machines ALWAYS worry about intermediate values. They have finite precision and cannot play like they don't.

OpenStudy (anonymous):

Just a guess from experience. 6 has no decimal point so it's assumed to be exact. 6*0.222*7.0 = 9.324 or 9.3 when 9.324 is rounded to the nearest tenth because of 7.0 .